OpenStudy (juana02):

Cable Cost The Schwimmer’s cost for cable television is given by c=75 + 25x , where x is the number of months of service. a) Graph c=75 + 25x , for x 50. b) From the graph of C, estimate the Schwimmer’s cost for 30 months of service. c) If the Schwimmer’s total cost for cable television is $975, estimate the number of months of service.

OpenStudy (danjs):

It is a linear equation, to graph it you need 2 points

OpenStudy (danjs):

i would use x=0 and x=50

OpenStudy (danjs):

c(0) = 75 + 25*0 = 75 (x , c(x)) = (0,75) is one point